# Photocathode characterisation for robust PICOSEC Micromegas precise-timing detectors

### v1

@@ Abstract
The PICOSEC Micromegas detector is a precise-timing gaseous detector based on a Cherenkov radiator coupled with a semi-transparent photocathode and a Micromegas amplifying structure, targeting a time resolution of tens of picoseconds for minimum ionising particles. Initial single-pad prototypes have demonstrated a time resolution below = 25 ps, prompting ongoing developments to adapt the concept for applications. The achieved performance is being transferred to robust multi-channel detector modules suitable for large-area detection systems requiring excellent timing precision. To enhance the robustness and stability of the PICOSEC Micromegas detector, research on robust carbon-based photocathodes, including Diamond-Like Carbon (DLC) and Boron Carbide (B4C), is pursued. Results from prototypes equipped with DLC and B4C photocathodes exhibited a time resolution of 32 ps and 34.5 ps, respectively. Efforts dedicated to improve detector robustness and stability enhance the feasibility of the PICOSEC Micromegas concept for large experiments, ensuring sustained performance while maintaining excellent timing precision.
@@ 1 Introduction
The intense interest in advancing technologies for precise-timing detectors has been driven by demanding environments anticipated in forthcoming High Energy Physics experiments. Meeting the criteria of achieving an excellent time resolution, ensuring stable long-term operation and providing large-area coverage is crucial for adapting the concept for physics applications. The PICOSEC Micromegas (hereafter PICOSEC) project [1] undertakes the efforts to develop a robust multi-channel gaseous detector, targeting a time resolution of ps for Minimum Ionising Particles (MIPs). Initial single-pad prototypes equipped with a Cesium Iodide (CsI) photocathode have demonstrated a time resolution below = 25 ps [1], prompting further developments [2] - [13]. Although preliminary measurements of alternative approaches have been conducted previously [2, 5, 11, 13], a comprehensive characterisation of different carbon-based photocathodes has not been reported. This work aims to advance the development of robust photocathodes for PICOSEC detectors while maintaining an excellent time resolution.
Within the scope of this paper, measurements of three different photocathode materials, including CsI, Diamond-Like Carbon (DLC) and Boron Carbide (B4C), are reported. The samples were manufactured by various research institutes, including the European Organization for Nuclear Research (CERN), the University of Science and Technology of China (USTC), the French Atomic Energy Commission (CEA) and the European Spallation Source (ESS).
@@ 2 PICOSEC Micromegas detection concept
The PICOSEC Micromegas detection concept is depicted in Fig. 1a [1]. The detector is designed to minimize time jitter from primary ionization in the gas by using a Cherenkov radiator. A charged particle passing through the crystal radiator generates a cone of ultraviolet (UV) photons which are partially converted into primary electrons on a photocathode coated on the window. Typically, the radiator in a PICOSEC detector is a 3 mm thick Magnesium Fluoride (MgF2) crystal which transmits light above 120 nm. The window is coated with a 3 nm thick chromium (Cr) conductive interfacial layer that serves as a contact for the high voltage (HV). The UV photon-to-electron converter is an 18 nm thick semi-transparent CsI photocathode known for its high Quantum Efficiency (QE) for photons below 200 nm. All primary electrons are created at the same surface, eliminating uncertainty about the ionisation location and thereby minimising time jitter. The gas volume is filled with a mixture of 80 Ne, 10 CF4 and 10 C2H6 at ambient pressure. The typical electric fields across the pre-amplification and amplification gaps are approximately 40 kV/cm and 20 kV/cm, correspondingly. In the presence of a strong electric field, the extracted electrons successively ionise gas molecules, causing electron multiplication, which first occurs in the pre-amplification gap and, after passing through the Micromegas mesh, continues in the amplification gap. The total gain achieved is on the level of . The amplified electrons move towards the anode, while the ions travel to the mesh. Their movement induce a signal on the anode, which is then amplified and read out by a digitiser. A typical PICOSEC waveform displaying a fast electron peak and a slow ion tail is illustrated in Fig. 1b. The leading edge of the electron peak determines a Signal Arrival Time (SAT).
Several optimisation studies were conducted on a 10 mm diameter active area single-pad PICOSEC detector to enhance high voltage stability, reduce noise, improve signal integrity and achieve uniform timing response, using a simplified assembly process [12]. The design includes improvements to the detector board, vessel, mechanical parts and electrical connections for both high voltage and signal. A new single-channel detector board is shown in Fig. 2.
@@ 3 Experimental methodology
R&D activities within the PICOSEC project have covered aspects ranging from design and production, through assembly, to measurements in both laboratory conditions as well as with particle beams. Prototypes were assembled in a clean-room environment and tested in the laboratory. The ASSET (A Small Sample Evaporation Test) setup was developed to characterise photocathodes, facilitating transparency and QE measurements, as well as ageing studies [14, 15]. The time resolution of the detectors was evaluated using 150 GeV/c muon beams at the CERN SPS H4 beamline.
@@ 3.1 ASSET photocathode characterisation device
The ASSET setup was developed to characterise photocathodes for gaseous radiation detectors, as in the PICOSEC project [5]. The primary aims are to quantify QE and transparency of the photocathodes as well as their degradation due to ion backflow. Fig. 3 provides an overview of the ASSET setup.
@@
Measurements utilise UV light from a system consisting of a deuterium lamp (McPherson, Model 632) and a Vacuum UltraViolet (VUV) monochromator (McPherson, Monarch). The system is flushed with high-purity nitrogen to prevent absorption of the short-wavelength portion of the UV spectrum. The monochromator is attached to the measurement chamber through an extension containing a collimating mirror chamber that focuses the light. A beam splitter divides the light into two beams. Two calibrated CsI photomultiplier tubes (PMTs) register the light intensity: one measures the amount of light at the sample position, while the other monitors light stability during measurements. The wavelength range for measurements is limited to about 140 nm on the low-wavelength side by the transparency of the MgF2 lens and to about 200 nm on the high-wavelength side by the deuterium lamp’s light intensity. The highest light intensity is observed at 160-180 nm.
@@
The ASSET setup allows for measurements in reflective and transmission modes. In both modes, the extraction grids record the current generated by electrons extracted from the photocathodes. The reflective mode measures the QE of photocathodes, while the transmission mode evaluates both QE and transparency. The measurements are conducted in a vacuum of 10-6 mbar, achieved using pre-pumps and turbomolecular pumps. Additionally, the device enables the the collection of determined amounts of ion current on samples to study potential charge-induced ageing effects. Further details about the ASSET photocathode characterisation device can be found in [5].
@@
The primary function of the ASSET setup is the measurement of photocurrent from photocathode samples. The QE is calculated from current measurements using the formula
@@
where is the number of electrons extracted from the sample (measured on the sample), is the number of photons that arrive at the sample (measured with the PMT), is the current measured on the sample (offset subtracted), is the current measured on the PMT (offset subtracted), is the calibration factor of PMT (depends on wavelength) and is the elementary charge.
The current measured by the PMT is used to calculate the transparency (T) of the samples, given by
@@
where is the PMT current measured with the sample in the measurement position and is the PMT current measured with the sample out of measurement position. Transparency measurements in ASSET was performed in the VUV range between 140 nm and 200 nm. Transparency can also be assessed across the visible light spectrum, from 200 nm to 800 nm, using a spectrophotometer (PerkinElmer, Lambda 650 UV/VIS Spectrometer).
@@
An important capability of the ASSET measurements is conducting ageing studies. The procedure starts with an X-ray beam entering the irradiation chamber filled with a gas mixture of Ar/CO2 (70/30 %) at ambient pressure that ionises the gas and creates primary electrons. The primary electrons are attracted to multiplication wires, where an avalanche occurs in the high electric field region near the wires, producing additional electrons and ions. The electrons from he avalanche are collected on the wires, while the ions move towards the grounded sample. This process accumulates charge on the sample, allowing for the quantification of the drop in QE after exposure.
@@ 3.2 Characterisation with particle beams
Particle beam campaigns are conducted to measure the time resolution of prototypes assembled in various configurations. The measurements are conducted using 150 GeV/c muon beams at the CERN SPS H4 beamline. The test setup includes a beam telescope facilitating triggering, timing and tracking capabilities. An example telescope configuration is shown in Fig. 4. Precise particle tracking is achieved utilising three triple Gas Electron Multiplier (GEM) detectors with a spatial resolution below 80 µm. These devices use APV25 front-end ASICs to shape the signals from all electrode channels, which are subsequently digitised using the Scalable Readout System (SRS) [16]. The GEMs are operated in a gas mixture of Ar/CO2 (70/30 %) at ambient pressure. A micro-channel plate photomultiplier tube (MCP-PMT, Hamamatsu, R3809U-50) serves as the timing reference and data acquisition (DAQ) trigger. The telescope can be utilised for testing several PICOSEC prototypes simultaneously.
@@
Different components individually contribute to the overall time resolution in a detector system, with the total time resolution being the sum of variances from each of these contributions. It can be represented by a simplified model
@@
The first term is the time resolution of the PICOSEC prototype measured with MIPs, where is the time resolution achieved with SPE events and is the number of photoelectrons extracted from the photocathode [2]. The second term is the contribution of the reference device. The third term is the time jitter caused by noise in the detector and the connected external electrical circuit. The time resolutions discussed in this paper include contributions from all these components.
The photocathode performance is assessed by comparing they generate during the process. The number is estimated as the ratio of the signal amplitude induced by a single photoelectron (SPE) to the signal amplitude induced by a MIP. SPE measurements are performed using a light-emitting diode (LED), while the multiple photoelectron (MPE) measurements are conducted with muon beams. The signals are amplified by a custom-developed charge-sensitive ARTH amplifier [2] and read out by an oscilloscope (LeCroy WR8104, 10 GS/s sampling rate, 1 GHz bandwidth).
The procedure for calculating involves several steps. Firstly, the maximum amplitude for each waveform is determined. Secondly, a histogram of all maximum amplitudes is plotted. The noise component is fitted using a Gaussian distribution, while the signal component is fitted using an equation based on a Pólya distribution of the form
where: is the number of charges produced, the mean avalanche size, and the shape parameter. From this Pólya distribution, the mean value is calculated. Finally, the mean amplitude of the MPE is divided by the mean amplitude of the SPE to obtain the NPE for a given photocathode. The is given by the formula
where is a MIP mean amplitude, is a SPE mean amplitude.
@@
Quantifying the time resolution of the PICOSEC detector requires a reference device with significantly superior timing precision. To fulfill this condition, the MCP-PMT with a time resolution below 6 ps in the central region [1, 2] serves as the reference for the PICOSEC prototypes. The devices are aligned to each other, ensuring that particles passed through the active area of both. To amplify and read out the signal, a custom-developed RF pulse amplifier cards optimized for PICOSEC (with built-in discharge protection up to 350 V, 650 MHz bandwidth, 38 dB gain, 75 mW power consumption [10, 17]) are used in conjunction with an oscilloscope.
@@
In the analysis, the constant fraction discrimination (CFD) method is employed to determine the timestamp at which the signal reaches a specific fraction of its peak amplitude, thereby addressing the issue of time walk [1]. The leading edge of the electron peak is fitted with a sigmoid function to identify the signal position in time at 20 of the its amplitude. The SAT is calculated as the difference between the timestamps of the PICOSEC detector and the reference device. The SAT distribution is analysed using a double Gaussian fit and time resolution of the detector system is computed as the standard deviation of this distribution.
@@
The results presented in this paper are derived from calculations conducted after applying several selection criteria (cuts) to the triggered events. Initially, a time window cut is implemented, selecting events within 300 ps of the median time difference of all recorded signals to exclude off-time events and noise fluctuations. Additional cuts are applied to set signal amplitude limits. Events with amplitudes below 1% of the dynamic range, categorised as empty events, and those above 99%, classified as saturated waveforms, are rejected. Furthermore, a geometrical cut is essential due to the Cherenkov radiator’s emission of a UV photon cone at approximately 45-degree angle. To ensure accurate time resolution measurements, only fully contained events within the detector’s active area should be included. Specifically, a cut of a 4 mm diameter circle around the pad center for a 10 mm diameter detector equipped with a 3 mm thick radiator (or a larger diameter, if otherwise specified) was implemented. Signals from tracks that pass outside this central region show reduced amplitude because of the partial loss of NPE beyond the area of the channel.

### v2

@@ Abstract
The PICOSEC Micromegas detector is a precise-timing gaseous detector based on a Cherenkov radiator coupled with a semi-transparent photocathode and a Micromegas amplifying structure, targeting a time resolution of tens of picoseconds for minimum ionising particles. Initial single-pad prototypes have demonstrated a time resolution below = 25 ps, prompting ongoing developments to adapt the concept for High Energy Physics applications, where sub-nanosecond precision is essential for event separation, improved track reconstruction and particle identification. The achieved performance is being transferred to robust multi-channel detector modules suitable for large-area detection systems requiring excellent timing precision. To enhance the robustness and stability of the PICOSEC Micromegas detector, research on robust carbon-based photocathodes, including Diamond-Like Carbon (DLC) and Boron Carbide (B4C), is pursued. Results from prototypes equipped with DLC and B4C photocathodes exhibited a time resolution of 32 ps and 34.5 ps, respectively. Efforts dedicated to improve detector robustness and stability enhance the feasibility of the PICOSEC Micromegas concept for large experiments, ensuring sustained performance while maintaining excellent timing precision.
@@ 1 Introduction
The intense interest in advancing technologies for precise-timing detectors has been driven by demanding environments anticipated in forthcoming High Energy Physics experiments [1]. Sub-nanosecond precision is crucial for distinguishing events occurring in rapid succession, improving track reconstruction and enabling particle identification through Time of Flight measurements. Meeting the criteria of achieving an excellent time resolution, ensuring stable long-term operation and providing large-area coverage is crucial for adapting the concept to be used in future experiments. The PICOSEC Micromegas (hereafter PICOSEC) project [2] undertakes the efforts to develop a robust multi-channel gaseous detector, targeting a time resolution of ps for Minimum Ionising Particles (MIPs). Initial single-pad prototypes equipped with Cesium Iodide (CsI) photocathodes have demonstrated a time resolution below = 25 ps [2], prompting further developments such as detector design optimisation [3], enhanced stability and robustness [4, 5, 6, 7], as well as scaling to larger areas [4, 8, 9, 10]. While the CsI photocathode exhibits high quantum efficiency (QE) and ultraviolet (UV) sensitivity, it is prone to damage from ion backflow and discharges, as well as humidity. Although preliminary measurements of alternative approaches have been conducted previously [4, 5, 6, 7], a comprehensive characterisation of different carbon-based photocathodes has not been reported. This work aims to advance the development of robust photocathodes for PICOSEC detectors while maintaining an excellent time resolution.
Within the scope of this paper, the characterisation of three different photocathode materials - CsI, Diamond-Like Carbon (DLC) and Boron Carbide (B4C) - and their performances as candidates for the PICOSEC detectors are reported. The samples were manufactured by various research institutes, including the European Organization for Nuclear Research (CERN), the University of Science and Technology of China (USTC), the French Atomic Energy Commission (CEA) and the European Spallation Source (ESS).
@@ 2 PICOSEC Micromegas detection concept
The PICOSEC Micromegas detection concept is detailed in [2] and depicted in Fig. 1a. The detector is designed to minimise time jitter from primary ionisation in the gas by using a Cherenkov radiator. A charged particle passing through the radiator generates a cone of UV photons which are partially converted into primary electrons on a photocathode coated on the crystal. Typically, the radiator in a PICOSEC detector is a 3 mm thick Magnesium Fluoride (MgF2) crystal which transmits light above 120 nm. The window is coated with a 3 nm thick chromium (Cr) conductive interfacial layer that serves as a contact for the high voltage (HV). The UV photon-to-electron converter is an 18 nm thick semi-transparent CsI photocathode known for its high QE for photons below 200 nm [2]. All primary electrons are created at the same surface, eliminating uncertainty about the ionisation location and thereby minimising time jitter.
The gas volume is filled with a mixture optimised for good time resolution by the COMPASS experiment [11], consisting of 80% Ne, along with 10% C2H6 and 10% CF4 to reduce electron diffusion, at ambient pressure. A stainless steel calendared woven mesh with 18 µm wire diameter, 45 µm opening, 30 µm thickness and about 50% optical transparency [2] serves a dual purpose: it protects the photocathode from ion feedback that could degrade its performance and ensures stability by preventing discharges. The pre-amplification gap between the cathode and the mesh is defined by a spacer made of two-layer copper-clad polyimide. The PICOSEC detector operates at HV levels close to, but not exceeding, the discharge limit. The typical electric fields across the pre-amplification and amplification gaps are around 30-40 kV/cm and 20 kV/cm, correspondingly. In the presence of a strong electric field, the extracted electrons successively ionise gas molecules, causing electron multiplication, which first occurs in the pre-amplification gap and, after passing through the mesh, continues in the amplification gap. The total gain achieved is on the level of .
The amplified electrons move towards the anode, while the ions travel to the mesh. Their movement induces a signal on the anode, which is then amplified and read out by a digitiser. A typical PICOSEC waveform displaying a fast electron peak and a slow ion tail is illustrated in Fig. 1b. The leading edge of the electron peak determines a Signal Arrival Time (SAT).
Several optimisation studies were conducted on the single-pad PICOSEC detector to enhance HV stability, reduce noise level, improve signal integrity and achieve uniform timing response, using a simplified assembly process. The design includes improvements to the detector board, vessel, mechanical parts and electrical connections for both HV and signal. A new single-pad detector board with a 10 mm diameter active area is shown in Fig. 2. A comprehensive overview of the mechanical design can be found in reference [3].
@@ 3 Experimental methodology
R&D activities within the PICOSEC project have covered aspects ranging from design and production, through assembly, to measurements in laboratory conditions as well as with particle beams. Prototypes were assembled in a clean-room environment and tested in the laboratory. The ASSET (A Small Sample Evaporation Test) setup was developed to characterise photocathodes, facilitating transparency and QE measurements, as well as ageing studies [5, 12, 13]. The time resolution of the detectors was evaluated using 150 GeV/c muon beams at the CERN SPS H4 beamline.
@@ 3.1 ASSET photocathode characterisation device
The ASSET setup was developed to characterise photocathodes for gaseous radiation detectors, as in the PICOSEC project [5]. The primary aims are to quantify QE and transparency of the photocathodes as well as their degradation due to ion backflow. Figure 3 provides an overview of the ASSET setup.
@@
Measurements utilise UV light from a system consisting of a deuterium lamp (McPherson, Model 632 [14]) and a Vacuum UltraViolet (VUV) monochromator (McPherson, Monarch [15]). The system is flushed with high-purity nitrogen to prevent absorption of the short-wavelength portion of the UV spectrum. The monochromator is attached to the measurement chamber through an extension containing a collimating mirror chamber that focuses the light. A beam splitter divides the light into two beams. Two calibrated CsI photomultiplier tubes (PMTs) register the light intensity: one measures the amount of light at the sample position, while the other monitors light stability during measurements. The wavelength range for measurements is limited to about 120 nm on the low-wavelength side by the transparency of the MgF2 lens and to about 200 nm on the high-wavelength side by the deuterium lamp’s light intensity. The highest light intensity is observed at 160-180 nm.
@@
The ASSET setup allows for measurements in reflective and transmission modes. The measurements are conducted in a vacuum of 10-6 mbar, achieved using pre-pumps and turbomolecular pumps. In both modes, the extraction grids record the current generated by electrons extracted from the photocathodes. The reflective mode measures the QE of photocathodes, while the transmission mode evaluates both QE and transparency.
@@
The main function of the ASSET setup is the measurement of current from photocathode samples. The QE is calculated from current measurements using the formula:
@@
where is the number of electrons extracted from the sample (measured on the sample), is the number of photons that arrive at the sample (measured on the PMT), is the current flowing from the sample to the wire grid, measured by a Keithley 6487 picoammeter [16], is the current measured on the PMT using the picoammeter, is the wavelength-dependent calibration factor for the PMT, as provided in the datasheet of the device and is the elementary charge.
The current measured by the PMT is used to calculate the transparency of the samples, given by:
@@
where is the PMT current measured with the sample in the measurement position, is the PMT current measured with the sample out of the position. Transparency measurements in ASSET was performed in the VUV range, usually between 120 nm and 200 nm. Transparency can also be assessed across the visible light spectrum, from 200 nm to 800 nm, using a spectrophotometer (PerkinElmer, Lambda 650 UV/VIS Spectrometer [17]).
@@
An important capability of the ASSET measurements is conducting ageing studies. The procedure starts with an X-ray beam entering the irradiation chamber filled with a gas mixture of Ar/CO2 (70/30 %) at ambient pressure that ionises the gas and creates primary electrons. The primary electrons are attracted to multiplication wires, where an avalanche occurs in the high electric field region near the wires, producing additional electrons and ions. The electrons from the avalanche are collected on the wires, while the ions move towards the grounded sample, where the charge is measured by the picoammeter [16]. This process accumulates charge on the sample, allowing for the quantification of the drop in QE after exposure. Further details about the ASSET photocathode characterisation device can be found in [5].
@@ 3.2 Characterisation with particle beams
Particle beam campaigns are conducted to measure the time resolution of prototypes assembled in various configurations. The measurements are conducted using 150 GeV/c muon beams at the CERN SPS H4 beamline. The test setup includes a beam telescope facilitating triggering, timing and tracking capabilities. An example telescope configuration is shown in Fig. 4. Precise particle tracking is achieved utilising three triple Gas Electron Multiplier (GEM) detectors with a spatial resolution below 80 µm. These devices use APV25 [18] front-end ASICs to shape the signals from all electrode channels, which are subsequently digitised using the Scalable Readout System (SRS) [19]. The GEMs are operated in a gas mixture of Ar/CO2 (70/30 %) at ambient pressure. A micro-channel plate photomultiplier tube (MCP-PMT, Hamamatsu, R3809U-50 [20]) serves as the timing reference and data acquisition (DAQ) trigger. The telescope can be utilised for testing several PICOSEC prototypes simultaneously.
@@
Different components individually contribute to the overall time resolution in a detector system, with the total time resolution being the sum of variances from each of these contributions. It can be represented by a simplified model:
The first term, , is the time resolution of the PICOSEC prototype measured with MIPs, including the detector’s noise coming from amplitude fluctuations. There is a dependence between the time resolution and the number of photoelectrons:
@@
where is the time resolution achieved with single photoelecton (SPE) events and is the number of photoelectrons extracted from a photocathode. For photocathodes with higher , a better time resolution is expected. The second term, , in Equation 3 is the contribution of the reference device. The third term, , is the time jitter caused by the connected external electrical circuit. The time resolutions discussed in this paper include contributions from all these components.
@@
Quantifying the time resolution of the PICOSEC detector requires a reference device with significantly superior timing precision. To fulfill this condition, the MCP-PMT with a time resolution below 6 ps in the central region [4] serves as the reference for the PICOSEC prototypes. The devices are aligned to each other, ensuring that particles passed through the active area of both. To amplify and read out the signal, a custom-developed RF pulse amplifier cards optimised for PICOSEC (with built-in discharge protection up to 350 V, 650 MHz bandwidth, 38 dB gain, 75 mW power consumption [10, 21]) and an oscilloscope (LeCroy WR8104, 10 GS/s sampling rate, 1 GHz bandwidth) are used.
@@
In the analysis, the constant fraction discrimination method is employed to determine the timestamp at which the signal reaches a specific fraction of its peak amplitude, thereby addressing the issue of time walk [2]. To reduce noise interference in the timing determination, the leading edge of the signal is fitted using a sigmoid function as it accurately represents its shape. The signal position in time at 20 of its amplitude is identified. The SAT values are computed as the difference between those timestamps, specifically by subtracting the MCP-PMT from the PICOSEC detector. The histogram of SAT values is fitted with a double Gaussian function, which effectively approximates the shape of the distribution and accounts for the variations between the SAT and the e-peak charge. The double Gaussian function is given by the equation:
where is a scaling factor, is a weighting parameter, is the mean of the distribution, is the standard deviation of the core Gaussian and is the standard deviation of the Gaussian describing the tail. The weighting parameter between the two Gaussian components enables the determination of the combined standard deviation of the overall distribution:
This combined standard deviation represents the time resolution of the detector system. Finally, the efficiency of the detector equipped with a given photocathode is calculated by dividing the number of events that induce a signal on the PICOSEC prototype by the number of events that induce a signal on the MCP-PMT when a charged particle passed through both devices.
@@
The results presented in this paper are derived from calculations conducted after applying several selection criteria (cuts) to the triggered events. Initially, a time window cut is implemented, selecting events within 300 ps of the median time difference of all recorded signals to exclude off-time events and noise fluctuations. Additional cuts are applied to set signal amplitude limits. Events with amplitudes below 1% of the dynamic range, categorised as empty events, and those above 99%, classified as saturated waveforms, are rejected. Furthermore, a geometrical cut is essential due to the Cherenkov radiator’s emission of a UV photon cone at approximately 45-degree angle. To ensure accurate time resolution measurements, only fully contained events within the detector’s active area should be included. Specifically, a cut of a 4 mm diameter circle around the pad center for a 10 mm diameter detector equipped with a 3 mm thick radiator (or a larger diameter, if otherwise specified) is implemented. Signals from tracks that pass outside this central region show reduced amplitude because of the partial loss of photoelectrons beyond the area of the channel. The efficiency of the detectors consistently remains above 95%, despite the implementation of corrections and cuts to the triggered events.
The photocathodes performance is additionally assessed by comparing they generate during the process. The number is estimated as a ratio of the signal amplitude induced by an SPE to the signal amplitude induced by a MIP for a detector at the same settings. SPE measurements are performed using a light-emitting diode (LED) to uniformly irradiate the photocathode and extract one photoelectron at a time, while the multiple photoelectron measurements are conducted with a muon beam. For the SPE measurements, an additional measurement with the LED off was performed to record only the noise. The signals are amplified by a custom-developed charge-sensitive ARTH amplifier [4] and read out by an oscilloscope.
The procedure for calculating involves several steps. Firstly, the maximum amplitude for each waveform is determined. Secondly, a histogram of all maximum amplitudes is plotted. For the LED-off measurement, the Gaussian distribution parameters are saved and later applied to the LED-on analysis. In the LED-on measurement, the noise component is fitted with a Gaussian distribution, while the signal component is fitted using a Pólya distribution of the form:
where: is the number of charges produced, is the mean avalanche size and is the shape parameter. In the muon beam measurement, only the Pólya distribution is used to fit the signal, since the noise component is negligible. The mean amplitude for the MIP, , is divided by the mean amplitude for the SPE, , to obtain the for a given photocathode:
Examples of histograms of the maximum amplitudes for SPE and MIP are presented in Fig. 5.